You are forecasting a stock to pay the following dividends:
$4.45 , $5.80 , $4.

The dividends will then begin declining at a rate of 8.5% for the foreseeable future. What is the intrinsic value of this stock if the required return is 14%?

I keep getting 18.42 but the answer shows 22.05 please advise.
Answer #1

Value after year 3=(D3*Growth rate)/(Required rate-Growth rate)

=(4*(1-0.085)/(0.14-(0.085)

=3.66/0.225

=16.267

Hence intrinsic value=Future dividend and value*Present value of discounting factor(rate%,time period)

=4.45/1.14+5.8/1.14^2+4/1.14^3+16.267/1.14^3

=$22.05(Approx).